The cost per hour of running an assembly line in a manufacturing plant is a function of the number of items produced per hour. The cost function is C(x)=0.3x^2-1.2x+2, where c(x) is the cost per hour in thousands of dollars, and x is the number of items produced per hour, in thousands. Determine the most economical production level.
Answer by stanbon(75887)

The cost per hour of running an assembly line in a manufacturing plant is a function of the number of items produced per hour. The cost function is C(x)=0.3x^2-1.2x+2, where c(x) is the cost per hour in thousands of dollars, and x is the number of items produced per hour, in thousands. Determine the most economical production level.
-------------
Find the minimum of C(x)
---------------
Minimum occurs when x = -b/2a = 1.2/(2*0.3) = 2
C(2) = 0.3(2)^2-1.2*2+2 = 0.80000
===============================
Cheers,
Stan H.